OpenStudy (anonymous):

Please help ill fan :D #1-Write an equation that will solve this problem: Two angles are supplementary. If the measure of the second angle is 3 times the sum of the measure of the first angle and 25, find the measures of the two angles. #2. Is x=-4 a solution of 4x+3(x+7)=-14 ? #3 -7h-17-6+4h=17-h #4 2(x+3)-6x-4=8+x PLEASE I need help PLEASE I will fan

OpenStudy (isaidavila):

The first a+b=180 b=3(a+25)

OpenStudy (isaidavila):

#2 no. -4 is not a solution of that equation. 4x+3x+21=-14 7x=-35 x=-5

OpenStudy (jmansedg):

this will help you http://openstudy.com/study#/updates/56bce5dce4b0598570438c37

OpenStudy (anonymous):

@isaidavila can u help with the other two to please?

OpenStudy (isaidavila):

#3 -3h-23=17-h -40=2h h=-20

OpenStudy (isaidavila):

#4 2x+6-6x-4=8+x -4x+2=8+x -5x=6 x=-6/5